# Benchmarking Cross-Domain Audio-Visual Deception Detection

### v1

@@ Abstract
Automated deception detection is crucial for assisting humans in accurately assessing truthfulness and identifying deceptive behavior. Conventional contact-based techniques, like polygraph devices, rely on physiological signals to determine the authenticity of an individual’s statements. Nevertheless, recent developments in automated deception detection have demonstrated that multimodal features derived from both audio and video modalities may outperform human observers on publicly available datasets. Despite these positive findings, the generalizability of existing audio-visual deception detection approaches across different scenarios remains largely unexplored. To close this gap, we present the first cross-domain audio-visual deception detection benchmark, that enables us to assess how well these methods generalize for use in real-world scenarios. We used widely adopted audio and visual features and different architectures for benchmarking, comparing single-to-single and multi-to-single domain generalization performance. To further exploit the impacts using data from multiple source domains for training, we investigate three types of domain sampling strategies, including domain-simultaneous, domain-alternating, and domain-by-domain for multi-to-single domain generalization evaluation. Furthermore, we proposed the Attention-Mixer fusion method to improve performance, and we believe that this new cross-domain benchmark will facilitate future research in audio-visual deception detection. Protocols and source code are available at https://github.com/Redaimao/cross_domain_DD.
@@ 1 Introduction
Traditional deception detection is often a contact-based method. It assesses whether someone is telling the truth or not by monitoring physiological responses like skin conductance and heart rate [12]. Experts’ Behavioral observation and analysis are another technique that evaluates changes in a person’s body language, speech patterns, and eye movements [13, 14]. However, such an assessment can be time-consuming and require significant expertise to perform accurately.
@@
To address the issue of cross-domain deception detection, we introduce a new benchmark that evaluates the generalization capacity of AI models using audio and visual features over publicly available datasets. Our benchmarking approach utilizes widely adopted audio and visual features, and we compare the single-to-single domain performance and multi-to-single domain generalization using different architectures. Specifically, for the multi-to-single setting, three domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, are implemented to conduct cross-domain testing. To further enhance performance, we propose an Attention-Mixer fusion method based on MLP-Mixer [19]. This benchmarking framework serves as an important tool for evaluating the effectiveness of audio-visual deception detection models in diverse contexts, which will help improve the capabilities of automated deception detection systems in real-world settings. Additionally, we hope our work will inspire further research on multimodal models that address domain shift issues. In summary, our main contributions include:
  • Introducing a new benchmark for evaluating the generalization capacity of AI models using audio and visual features across different domains.
  • Comparing the single-to-single domain and multi-to-single domain generalization using different architectures.
  • Providing three domain sampling strategies i.e., domain simultaneous, domain alternating, and domain-by-domain, to conduct multi-to-single cross-domain testing.
  • Proposing the Attention-Mixer fusion method to enhance performance.
@@ 2.1 Cues to Deception
Levine et al. [24] reviewed the status quo and provided a new perspective on the theories of deception. They pointed out that lying usually happens when problematic information is involved. It is critical to understand the verbal content in the context. Vrij et al. [25] realized that interviewers play a vital role in eliciting and enhancing cues to deceit. The authors proposed the “interviewing to detect deception” technique to open a new path in the deception detection research field. They argued that different psychological states can be exploited by adopting appropriate interview techniques of liars and truth-tellers. Hirschberg et al. [26] proposed a method to distinguish deceptive from non-deceptive speeches by a large corpus. They also conducted experiments using acoustic, lexical, and speaker-dependent features, which showed improved performance by combining multiple feature sets. Warren et al. [27] conducted experiments to investigate the relationship between affective facial expressions to deception. The results indicated that leaked emotions with the incongruous intended message can provide useful cues to deception, which supported the nonverbal leakage theory [28, 29].
@@ 2.2 Multimodal Deception Detection
Recent works for deception detection usually use verbal and non-verbal features and propose effective fusion methods [30, 15, 16, 17]. For example, some works utilized facial features from RGB images to perform deception detection [15, 17, 5]. To capture facial movements, facial action units (AUs) were utilized. Other features, such as facial expression, were also adopted [6, 30]. Besides visual features, many works incorporated audio features to boost performance [7, 6, 15]. For example, Wu et al. [6] used MFCC (Mel-frequency Cepstral Coefficients) features and Karimi et al. [15] used raw audio. Most of the recent works mentioned above have considered multimodal fusion approaches that extract visual, audio, and text information to boost performance. In addtion to visual, audio and text, Karnati et al. [16] exploited physiological signals, i.e., EEG representations for deception detection.
To better fuse the multimodal features, different fusion methods were proposed, which can be broadly categorized into feature-level fusion and decision-level fusion. Specifically, feature-level fusion focused on producing better multimodal embeddings and used the linear layers to extract crossmodal dynamics [15, 17, 5, 6, 7]. Rather than that, decision-level fusion aimed to fuse multimodal dynamics at a late stage, to reduce computational complexity and learn good marginal representations [16, 7].
However, previous works on multimodal deception detection did not consider cross-domain issues that occur from one domain to another, which is the focus of this work.
@@ 3.1 Audio and Visual Feature Learning
$\mathcal{L}=\frac{1}{N}\sum_{i=i}^{N}\left(\sum_{m={f,b,a}}\mathcal{L}_{m,i}+\lambda\mathcal{L}_{g,i}\right),$ (3)
where $N$ is the number of data samples and $\lambda$ is a trade-off parameter between modality loss and fusion loss. $\lambda$ is set to 0.5 in our experiments.
@@ 3.2 Cross-domain Generalization
where $X=(X_{f},X_{b},X_{a})$, $Y$ is the label, $\mathcal{L}$ is the loss function, and $\mathbb{E}$ is the expectation.
When $M=1$, it is a Single-to-single Cross-domain Generalization task, where the modal is trained on one training dataset and tested on another dataset. When $M\geqslant 2$, we propose three strategies to learn from multiple domains for the Multi-to-single Cross-domain Generalization. Let $B$ denote one batch of training data with a size of $N_{B}$. Given multiple training domains $\mathcal{S}_{train}=\{S_{j}|j=1,\cdots,M\}$, $B$ is a set of training data sampled from $\mathcal{S}_{train}$.
Domain-Simultaneous means to train multiple domains in parallel within each batch of data. In domain simultaneous training, the $k-$th batch of training data is a group of samples from different domains, i.e., $B^{k}=(b_{S_{1}}^{k},\cdots,b_{S_{M}}^{k})$, $k\in[1,\cdots K]$, where $b_{S_{j}}^{k}$ is the batch samples from domain $S_{j}$ for $j=1,\cdots,M$, $K$ is the number of batches during training. The total number of $b_{S_{j}}^{k}$ is $N_{B}$. As shown in Fig. 2 (b), each training batch contains smaller batch samples from all the source domains during training. Models are trained to learn from different domains simultaneously by feeding the mixed batch data.
@@ 3.3 Attention-Mixer Fusion
Besides investigating cross-domain sampling strategies, inspired by [19], we propose Attention-Mixer Fusion to enhance the performance by fusing audio-visual modalities, where the attention mixer layer takes multimodal features as input to produce fused features. In particular, an attention mixer layer is composed of unimodal MLP layers, self-attention layers [37], and crossmodal MLP layers. First, for batch size $N_{B}$, the input features from different modalities are concatenated and projected to be a tensor $U\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ by a Liner layer, followed by several attention mixer layers, where $N_{m}$ is the number of input modalities. Specifically, the unimodal MLP layer, the self-attention layer, and the crossmodal MLP layer can be respectively described as
$\small U^{*,*,i}=F_{g}^{*,*,i}+\mathbf{W_{2}}\ \sigma(\mathbf{W_{1}}\ LN(F_{g}^{*,*,i})),\ i=\left[1,D\right],$ (5)
$\footnotesize U=\left[\left(softmax\left(U{\mathbf{W_{3}}(U\mathbf{W_{4}})^{T}}\over{\sqrt{D}}\right)U\mathbf{W_{5}}\right)_{h}\right]\mathbf{W_{6}},\ h=[1,H],$ (6)
$\small U^{*,j,*}=U^{*,j,*}+\mathbf{W_{8}}\ \sigma(\mathbf{W_{7}}\ LN(U^{*,j,*})),\ j=\left[1,N_{m}\right],$ (7)
@@
where $LN(\cdot)$ denotes the Layer Normalization, $\mathbf{W}_{1-8}$ are trainable weights, $H$ is the number of heads in multihead self-attention, and * denotes all the entries in that dimension. Several attention mixer layers are stacked as a deep block, which is set as a hyperparameter in practice. We set it to 6 in our experiment. Finally, $U\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ is reduced to $U\in\mathbb{R}^{N_{B}\times N_{m}\times 1}$ by obtaining the mean value on the feature dimension. In Eq. 5, the unimodal MLP layer is conducted along the feature dimension to learn the dynamics in each unimodal feature. Eq. 6 shows the multi-head self-attention operation on the tensor $U$, which further explores the attention between the unimodal features. In Eq. 7, the crossmodal MLP layer learns the dynamics across the modality dimension from the corresponding feature tokens.

### v2

@@ Abstract
Automated deception detection is crucial for assisting humans in accurately assessing truthfulness and identifying deceptive behavior. Conventional contact-based techniques, like polygraph devices, rely on physiological signals to determine the authenticity of an individual’s statements. Nevertheless, recent developments in automated deception detection have demonstrated that multimodal features derived from both audio and video modalities may outperform human observers on publicly available datasets. Despite these positive findings, the generalizability of existing audio-visual deception detection approaches across different scenarios remains largely unexplored. To close this gap, we present the first cross-domain audio-visual deception detection benchmark, that enables us to assess how well these methods generalize for use in real-world scenarios. We used widely adopted audio and visual features and different architectures for benchmarking, comparing single-to-single and multi-to-single domain generalization performance. To further exploit the impacts using data from multiple source domains for training, we investigate three types of domain sampling strategies, including domain-simultaneous, domain-alternating, and domain-by-domain for multi-to-single domain generalization evaluation. We also propose an algorithm to enhance the generalization performance by maximizing the gradient inner products between modality encoders, named “MM-IDGM”. Furthermore, we proposed the Attention-Mixer fusion method to improve performance, and we believe that this new cross-domain benchmark will facilitate future research in audio-visual deception detection. Protocols and source code are available at https://github.com/Redaimao/cross_domain_DD.
@@ 1 Introduction
Traditional deception detection is often a contact-based method. It assesses whether someone is telling the truth or not by monitoring physiological responses like skin conductance and heart rate [12, 13, 14]. Experts’ Behavioral observation and analysis are another technique that evaluates changes in a person’s body language, speech patterns, and eye movements [15, 16]. However, such an assessment can be time-consuming and require significant expertise to perform accurately.
@@
To address the issue of cross-domain deception detection, we introduce a new benchmark that evaluates the generalization capacity of AI models using audio and visual features over publicly available datasets. Our benchmarking approach utilizes widely adopted audio and visual features, and we compare the single-to-single domain performance and multi-to-single domain generalization using different architectures. Specifically, for the multi-to-single setting, three domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, are implemented to conduct cross-domain testing. To further enhance performance, we propose Multimodal Inter-Domain Gradient Matching (MM-IDGM) to improve multi-to-single domain generalization performance and an Attention-Mixer fusion method based on MLP-Mixer [21]. In our experiments, we compared single-to-single and multi-to-single domain generalization, incorporating three sampling strategies, one domain generalization method, and six fusion methods across two modalities. The benchmarking results among different features reveal that the best average fusion accuracy was 58.58% for single-to-single cross-domain generalization and 59.13% for multi-to-single cross-domain generalization. Moreover, for MM-IDGM, we achieved the highest average accuracy of 59.69% based on domain-simultaneous sampling. This benchmarking framework serves as an important tool for evaluating the effectiveness of audio-visual deception detection models in diverse contexts, which will help improve the capabilities of automated deception detection systems in real-world settings. Additionally, we hope our work will inspire further research on multimodal models that address domain shift issues. In summary, our main contributions include:
  • Introducing a new benchmark for cross-domain audio-visual deception detection: We present a comprehensive benchmark that evaluates the generalization capacity of AI models using audio and visual modalities across multiple domains in the deception detection task.
  • Comparing single-to-single and multi-to-single domain generalization across different architectures: We conduct extensive experiments comparing the performance of AI models in single-to-single and multi-to-single domain generalization over different architectures and modalities inputs.
  • Testing three domain sampling strategies for multi-to-single domain generalization: We propose and evaluate three distinct domain sampling strategies, i.e., domain simultaneous, domain alternating, and domain-by-domain, to assess their effectiveness and offer diverse approaches to handling domain discrepancies.
  • Introducing Multimodal Inter-Domain Gradient Matching (MM-IDGM) for domain generalization: We propose MM-IDGM to improve multi-to-single domain generalization by aligning gradient updates between modalities, enhancing the model’s ability to transfer knowledge across different domains.
  • Proposing the Attention-Mixer fusion method for enhanced multimodal performance: We introduce the Attention-Mixer fusion method that combines MLP-Mixer and self-attention layers to capture both intra- and inter-modal interactions better, leading to improved fusion performances in audio-visual deception detection tasks.
@@ 2.1 Cues to Deception
Levine et al. [26] reviewed the status quo and provided a new perspective on the theories of deception. They pointed out that lying usually happens when problematic information is involved. It is critical to understand the verbal content in the context. Vrij et al. [27] realized that interviewers play a vital role in eliciting and enhancing cues to deceit. The authors proposed the “interviewing to detect deception” technique to open a new path in the deception detection research field. They argued that different psychological states can be exploited by adopting appropriate interview techniques of liars and truth-tellers. Warren et al. [28] conducted experiments to investigate the relationship between affective facial expressions to deception. The results indicated that leaked emotions with the incongruous intended message can provide useful cues to deception, which supported the nonverbal leakage theory [29].
@@ 2.2 Multimodal Deception Detection
Recent advances in deception detection have increasingly incorporated both verbal and non-verbal cues, leveraging multimodal data to enhance detection performance. Often, those methods integrate features from different modalities, such as visual, audio, and textual modalities, and propose fusion strategies to combine them efficiently [30, 17, 18, 19]. In this section, we review and discuss the strengths and limitations of recent audio-visual deception detection methods, including the fusion techniques used and the gaps in addressing cross-domain issues.
Visual Features in Deception Detection. Many recent works have utilized facial features extracted from RGB images to perform deception detection [17, 19, 5]. They focus on capturing facial movements and expressions, which are crucial indicators of deception. Many studies have employed Facial Action Units (AUs) as a key feature [17, 19] to analyze subtle facial movements. AUs provide detailed representations of specific facial muscle activations, making them highly useful for identifying deceptive cues. Additionally, gaze movements and facial expressions have been widely adopted as visual features in deception detection models [6, 30, 31, 32]. By capturing emotional states that may signal deception, these methods can further enhance the detection capabilities.
Audio Features for Deception Detection. Along with visual cues, many recent works have incorporated audio features to boost the performance of deception detection systems. Audio signals can provide complementary information, such as changes in speech patterns, tone, and hesitation, which are often associated with deceptive behavior [7, 6, 17]. For example, Wu et al. [6] used MFCC (Mel-frequency Cepstral Coefficients) features, and Karimi et al. [17] explored the use of raw audio signals for deception detection, demonstrating the potential of unprocessed data to capture subtle vocal cues indicative of deception.
Multimodal Fusion Techniques. The integration of multiple modalities, such as visual, audio, and textual information, has been shown to improve deception detection performance significantly. Most recent works mentioned above have considered multimodal fusion approaches that extract visual, audio, and text information to boost performance [33, 34]. In addition to visual, audio, and text, Karnati et al. [18] exploited physiological signals, i.e., EEG representations for deception detection. Fusion methods can be broadly categorized into two types: feature-level fusion and decision-level fusion. Feature-level fusion combines raw or intermediate features from each modality into a single representation, allowing the model to learn cross-modal dynamics [17, 19, 5, 6, 7]. This method often uses different encoders, linear layers, or fusion modules to blend features and capture interactions between modalities. On the other hand, decision-level fusion merges the outputs or produced logits of individual unimodal models at a later stage. It can reduce computational complexity and enable the system to learn effective marginal representations from each modality [18, 7]. Both approaches have shown promising results in improving the performance of deception detection systems.
Cross-Domain Generalization Despite significant progress, there still exists the challenge of cross-domain generalization in multimodal deception detection. Previous works have mainly focused on optimizing unimodal or fusion methods within a single domain, without considering the domain shift when the system is applied to different environments or populations. For example, models trained on controlled lab data may not generalize well to real-world settings, where the recording conditions and communication styles may introduce significant variability.
In this work, we aim to build a benchmark for cross-domain generalization performance on the widely used audio-visual deception detection datasets, which is crucial for evaluating and improving the robustness of deception detection models in different scenarios. Establishing such a benchmark will provide clearer comparisons, highlight model weaknesses, and guide the development of systems across different domains.
@@ 3.1 Audio and Visual Feature Learning
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left(\sum_{m={f,b,a}}\mathcal{L}_{m,i}+\lambda\mathcal{L}_{g,i}\right),$ (3)
where $N$ is the number of data samples and $\lambda$ is a trade-off parameter between modality loss and fusion loss. $\lambda$ is set to 0.5 in our experiments. For score fusion, the fusion score is the average of the individual predicted scores for each class, and the cross-entropy loss function is applied to the fused score. For feature fusion, individual features are the inputs for the fusion block.
@@ 3.2 Cross-domain Generalization
where $X=(X_{f},X_{b},X_{a})$ is the input, $Y$ is the label, $\mathcal{L}$ is the loss function, and $\mathbb{E}$ denotes the expectation.
When $M=1$, it is a Single-to-single Cross-domain Generalization task, where the modal is trained on one training dataset and tested on another dataset. When $M\geqslant 2$, we propose three strategies to learn from multiple domains for the Multi-to-single Cross-domain Generalization. Previous works [41, 42] have indicated that different sampling strategies for domain generalization may result in different performances. Therefore, in this work, we propose three multi-to-single domain generalization sampling methods and compare their performance. Let $B$ denote one batch of training data with a size of $N_{B}$. Given multiple training domains $\mathcal{S}_{train}=\{S_{j}|j=1,\cdots,M\}$, $B$ is a set of training data sampled from $\mathcal{S}_{train}$.
Domain-Simultaneous means to train multiple domains in parallel within each batch of data. In domain simultaneous training, the $k-$th batch of training data is a group of samples from different domains, i.e., $B^{k}=(b_{S_{1}}^{k},\cdots,b_{S_{M}}^{k})$, $k\in[1,\cdots K]$, where $b_{S_{j}}^{k}$ is the batch samples from domain $S_{j}$ for $j=1,\cdots,M$ and $K$ is the number of batches during training. The total number of $b_{S_{j}}^{k}$ is $N_{B}$. As shown in Fig. 2 (b), each training batch contains smaller batch samples from all the source domains during training. Models are trained to learn from different domains simultaneously by feeding the mixed batch data. This approach helps the model learn domain-invariant features by exposing it to data from multiple domains simultaneously. However, it may cause the model to focus on “average” features across domains, leading to poor generalization to the unseen domains. Mixing domains ensures broader feature exposure, but can dilute domain-specific nuances that are useful for recognizing patterns in unseen domains.
@@
Multimodal Inter-Domain Gradient Matching. Beyond the benchmarking on the three different multi-to-single sampling strategies, inspired by Shi et al. [43], we propose Multimodal Inter-Domain Gradient Matching (MM-IDGM). IDGM was proposed by Shi et al. [43] and uses an effective optimization algorithm named “Fish” to approximate the second-order derivative in IDGM and to reduce computational cost. The main idea of IDGM is to maximize the gradient inner product to align the gradient directions across domains, so as to learn the weights that can produce the closer input-output correspondence. However, considering the multimodal nature of our task, directly optimizing IDGM in multimodal models may not be desirable. Different from IDGM in [43] which aims to learn the features that are invariant across domains, MM-IDGM prioritizes learning the invariant features across domains for the same modalities by maximizing the gradient inner product between modality encoders. These inner gradient products are then dynamically adjusted based on the unimodal losses from previous steps. The remaining parts of the models are then optimized across domain by IDGM. The MM-IDGM algorithm is developed based on the “Fish” algorithm and is presented in Algorithm 1.
Algorithm 1 MM-IDGM for Modalities $m_{1}$ and $m_{2}$
$L_{m_{1}}$, $L_{m_{2}}$triangleright from the previous step
while  iterations $\leq$ maximum iterations do
     $\tilde{\theta}_{m_{1}}\leftarrow\theta_{m_{1}}$, $\tilde{\theta}_{m_{2}}\leftarrow\theta_{m_{2}}$, $\tilde{\theta}_{r}\leftarrow\theta_{r}$
     for $S_{i}\in$permute$\{S_{1},S_{2},\cdots,S_{M}\}$ do
         Sample batch $s_{i}\sim S_{i}$
         $g_{m_{1},i}=\mathbb{E}_{s_{i}}[\frac{\partial L(X_{m_{1}},Y);\tilde{\theta}_{m_{1}}}{\partial\tilde{\theta}_{m_{1}}}]$triangleright encoder $m_{1}$
         $g_{m_{2},i}=\mathbb{E}_{s_{i}}[\frac{\partial L(X_{m_{2}},Y);\tilde{\theta}_{m_{2}}}{\partial\tilde{\theta}_{m_{2}}}]$triangleright encoder $m_{2}$
         $g_{r,i}=\mathbb{E}_{s_{i}}[\frac{\partial L(X_{r},Y);\tilde{\theta}_{r}}{\partial\tilde{\theta}_{r}}]$triangleright other
         Update $\tilde{\theta}_{m_{1}}\leftarrow\tilde{\theta}_{m_{1}}-\alpha(\frac{L_{m_{2}}}{L_{m_{1}}+L_{m_{2}}})g_{m_{1},i}$
         Update $\tilde{\theta}_{m_{2}}\leftarrow\tilde{\theta}_{m_{2}}-\alpha(\frac{L_{m_{1}}}{L_{m_{1}}+L_{m_{2}}})g_{m_{2},i}$
         Update $\tilde{\theta}_{r}\leftarrow\tilde{\theta}_{r}-\alpha g_{r,i}$
     end for
     Update $\theta\leftarrow\theta+\epsilon(\tilde{\theta}-\theta))$
     Update $L_{m_{1}}$, $L_{m_{2}}$
end while
MM-IDGM first performs inner gradient updates on a cloned version of the original model, denoted as $\tilde{\theta}$, focusing on each modality encoder individually. The gradients for each modality encoder are dynamically adjusted based on previous unimodal losses, ensuring a more balanced learning rate across modalities. Finally, the weights of the original model, $\theta$, are updated by applying a weighted difference between the cloned model and the original one.
@@ 3.3 Attention-Mixer Fusion
Besides investigating cross-domain sampling strategies, inspired by [21], we propose Attention-Mixer Fusion to enhance the performance by fusing audio-visual modalities, where the attention mixer layer takes multimodal features as input to produce fused features. In particular, an attention mixer layer is composed of unimodal MLP layers, self-attention layers [44], and crossmodal MLP layers. First, for batch size $N_{B}$, the input features from different modalities are concatenated and projected to be a tensor $F_{g}\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ by a Liner layer, followed by several attention mixer layers, where $N_{m}$ is the number of input modalities. Specifically, the unimodal MLP layer, the self-attention layer, and the crossmodal MLP layer can be respectively described as
$U^{*,*,i}=F_{g}^{*,*,i}+\mathbf{W_{2}}\ \sigma(\mathbf{W_{1}}\ LN(F_{g}^{*,*,i})),\ i=\left[1,D\right],$ (5)
$\small U=\left[\left(softmax\left(U{\mathbf{W_{3}}(U\mathbf{W_{4}})^{T}}\over{\sqrt{D}}\right)U\mathbf{W_{5}}\right)_{h}\right]\mathbf{W_{6}},\ h=[1,H],$ (6)
$U^{*,j,*}=U^{*,j,*}+\mathbf{W_{8}}\ \sigma(\mathbf{W_{7}}\ LN(U^{*,j,*})),\ j=\left[1,N_{m}\right],$ (7)
@@
where $LN(\cdot)$ denotes the Layer Normalization, $\mathbf{W}_{1-8}$ are trainable weights, $H$ is the number of heads in multihead self-attention, and * denotes all the entries in that dimension. Several attention mixer layers are stacked as a deep block, which is set as a hyperparameter in practice. We set it to 6 in our experiment. Finally, the output tensor $U\in\mathbb{R}^{N_{B}\times N_{m}\times D}$ is reduced to $U\in\mathbb{R}^{N_{B}\times N_{m}\times 1}$ by obtaining the mean value on the feature dimension. In Eq. 5, the unimodal MLP layer is conducted along the feature dimension to learn the dynamics in each unimodal feature. Eq. 6 shows the multi-head self-attention operation on the tensor $U$, which further explores the attention between the unimodal features. In Eq. 7, the crossmodal MLP layer learns the dynamics across the modality dimension from the corresponding feature tokens. In contrast to the MLP-Mixer, which utilizes a simple two-layer MLP structure, our approach introduces a self-attention layer between the two MLP layers. This allows for more efficient global operation from MLP-Mixer and also adaptively weighing token interaction by self-attention. This architectural enhancement allows our model to more effectively capture both intra- and inter-modal interactions, making it especially well-suited for processing multimodal data such as audio and visual inputs.
4 Experiments
In this part, extensive experiments are conducted to benchmark the cross-domain performances of public deception detection datasets.
4.1 Databases and Metrics
4.2 Implementation Details
Model Selection. Models for audio and visual modalities were selected to fit the data volume. For face frames, we adopted ResNet18 [35] and Gate Recurrent Unit (GRU) [46] models for facial feature extraction and temporal modeling, respectively. Two-layer multilayer perception (MLP) [47] models were used for AUs, gaze, and affect feature representation. Specifically, the first MLP Layer initially operates on the dimension that represents the number of features, with sizes following the sequence: “input dimension $\rightarrow$ 64 $\rightarrow$ 128 $\rightarrow$ 1. Afterward, it reshapes the tensor, and the MLP Layer processes on the feature dimension, where the sizes progress as “64 $\rightarrow$ 32 $\rightarrow$ 16 $\rightarrow$ 2. The MLP networks have identical structures, except for their first linear layers, as the input dimensions differ. For the audio-based Mel spectrogram, we used the ResNet18 [35] model for time-frequency feature representation. For audio waveforms, the Wave2Vec [48] model was applied for audio feature extraction.